Find the cross product of (3, 2, 0) and (2, −1, −2), and use the dot product to confirm that your answer is correct. 


1
Expert's answer
2021-05-02T06:42:54-0400

To find the cross product, we form a determinant the first row of which is a unit vector, the second row is our first vector, and the third row is our second vector

"\begin{pmatrix}
 3 \\
 2 \\
 0
\end{pmatrix} \times \begin{pmatrix}
 2 \\
 -1 \\
 -2
\end{pmatrix}= \begin{vmatrix}
 i & j & k\\
 3 & 2 & 0\\
 2 & -1 & -2
\end{vmatrix}"


"\begin{pmatrix}
 3 \\
 2 \\
 0
\end{pmatrix} \times \begin{pmatrix}
 2 \\
 -1 \\
 -2
\end{pmatrix}= \begin{vmatrix}
 2 & 0 \\
 -1 & -2 
\end{vmatrix}i - \begin{vmatrix}
 3 & 0 \\
 2 & -2 
\end{vmatrix} j+\begin{vmatrix}
 3 & 2 \\
 2 & -1
\end{vmatrix}k"


"((2).(-2)-(0).(-1))i-((3).(-2)-(0).(2))j+((3).(-1)-(2).(2))k"


"-4i+6j-7k"

"\begin{pmatrix}
 3 \\
 2 \\
 0
\end{pmatrix} \times \begin{pmatrix}
 2 \\
 -1 \\
 -2
\end{pmatrix}= \begin{pmatrix}
 -4 \\
 6 \\
 -7
\end{pmatrix}"


To confirm that this vector is perpendicular we can check the dot product is zero:

"\u27e8 3, 2 , 0\u27e9\u22c5\u27e8 -4 , 6, -7\u27e9 = -12+ 12 + 0 = 0"

"\u27e8 2, -1 , -2\u27e9\u22c5\u27e8 -4 , 6, -7\u27e9 = -8 - 6 + 14= 0"